# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

### v1

@@ 4. The frequency of any value $x\in[d]$ can be estimated as
The only output that depends on the value of $X^{(i)}$ is $z^{(i)}$, which is perturbed by an $\epsilon$-LDP mechanism. The postprocessing of the outputs from an LDP mechanism does not change its privacy guarantee [16], so we have proved:
@@ 3.1 Expectation and its Bias
$E[\hat{f}(x)]\\ =\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1}$ (3)
@@
If and only if $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$,
@@
The proof of Theorem 3.2 and Corollary 3.1 is presented in Appendix A. It is notable that $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is equivalent to $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, where $h\in\mathcal{H}$ represents uniform sampling, which is also the definition of a pairwise-independent hashing family. Thus, Corollary 3.1 can be rephrased as:
@@
The pairwise independence requires the size of $\mathcal{H}$ to be $d^{2}$, so the communication cost of sending one hash function becomes $2\log_{2}(d)$. The creation of such a hashing family is presented in Section 3.5.
@@ 3.2 Variance
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]},$ (5)
@@
The lemma, theorem, and corollary of this section are proven in Appendix B.
3.3 Accuracy Analysis
The mean squared error (MSE) measures the accuracy of $\hat{f}(x)$, which is defined as
@@ 3.3 Accuracy Analysis
$MSE(g(x))=\frac{1}{n(1-\bar{c}(x,x^{\prime}))^{2}}\biggl{[}\\ (1-f(x))\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (10)
@@
Note that Theorem 10 is not applicable to the original CMS [7] because its $\bar{c}(x,x^{\prime})$ is not constant. We will analyze its accuracy in Section 3.6.
Besides MSE, one may also use the confidence interval (CI) of $\hat{f}(x)$ to measure accuracy. If $\hat{f}(x)$ is unbiased, its CI is the same as its concentration bound. Moreover, $\hat{f}(x)$ is equivalent to the sum of bounded random variables (i.e., $\sum\hat{y}$), so its bounds can be derived from the Chernoff bound. Below is the concentration bound for $\hat{f}(x)$ from the unbiased CMS using randomized response (referred to as CMS+RR):
@@
This theorem is proven in Appendix C. Note that CMS+RR will be proven to be optimized in Section 3.4, so this theorem is applicable directly to the optimized CMS.
3.4 Optimized CMS
This section explores the optimized parameters of CMS to reduce the $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE defined in Section 2.2. It is notable that $l_{2}$ loss is connected to MSE as follows:
@@ 3.4 Optimized CMS
We start with the LDP process $R$. [9] summarized the state-of-the-art frequency estimation algorithms at that time. When considering different privacy regimes and original frequencies, i.e., $f(x)$, we found that randomized response (RR), symmetric RAPPOR (s-RP), and asymmetric RAPPOR (a-RP) are optimized in some privacy regimes. Though more optimized algorithms exist, their performance is either as optimized as these algorithms in their respective regimes, or they can be converted to CMS+RR (as shown in Section 4). Thus, we only consider RR, s-RP, and a-RP as candidates for the LDP process $R$.
The variance of the these mechanisms are listed below:
$Var(RR(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}e^{\epsilon}(m-1)&\text{ if }a=b\\ e^{\epsilon}+m-2&\text{ if }a\neq b,\end{cases}$ (12)
and
Then, we explore the optimized $m$ for different LDP processes and compare their loss functions and the worst-case MSE. We start with the worst-case MSE. Here, we allow an analyst to have prior knowledge of the upper bound of $f(x)$, and we derive:
@@
If the upper bound of $f(x)$ is known as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$, where $\mathbf{x}$ are the values of interest, the worst-case MSE among $\mathbf{x}$ is optimized as
@@
which can be achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to
@@
The theorem is proven in Appendix D. Notably, $f^{*}$ is prior knowledge of an analyst and could be larger than $\max_{x}f(x)$, which is unknown to the analyst. Subsequently, we have:
@@
If the upper bound of $f(x)$ is unknown, i.e., $f^{*}=1$, the worst-case MSE is optimized as
$\widehat{MSE}(\hat{f})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}},$
which can be achieved when $R$ is randomized response, and $m$ is the closest integer to $1+e^{\epsilon/2}$
Hereinafter, we assume no prior knowledge of $f(x)$ unless explicitly mentioned, i.e., $f^{*}=1$ by default. On the other hand, if an analyst is interested in some rare values, and the corresponding $f^{*}$ is around zero, we can substitute $f^{*}=0$ into Theorem 15 and obtain $m=1+e^{\epsilon}$, which should be sufficient to approximate the optimized $m$. Its worst-case variance is derived as:
$\widehat{MSE}(\hat{f})=\frac{1}{n}[f^{*}+\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}].$ (16)
We can see $m=1+e^{\epsilon}$ works well if $f^{*}$ is not a dominant term in Eq. (16).
Next, we derive the optimized $l_{1}$ and $l_{2}$ losses:
Theorem 3.7.
The $l_{1}$ and $l_{2}$ losses are optimized as
@@
$l_{2}^{*}(\hat{f})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}},$
respectively, where
$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}.$
Both optimalities are achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$.
Corollary 3.5.
When the dictionary size $d$ is large enough, the optimized $l_{1}$ and $l_{2}$ losses approach
@@
The theorem and corollary above are proven in Appendix D. It should be noted that the CMS using symmetric or asymmetric RAPPOR as its LDP process cannot serve as the optimized CMS for the following reasons: (1) their performance in reducing worst-case MSE or $l_{1}$ / $l_{2}$ losses is outperformed by CMS+RR in some conditions, and (2) they have very high communication costs. Please refer to Appendix D for more details.
One caveat of the setting to minimize $l_{1}$ / $l_{2}$ losses is that it will lead to a worst-case MSE larger than the optimized value. Assuming $d$ is large enough, and $m=1+e^{\epsilon}$, we have
$\widehat{MSE}(\hat{f})=Var(\hat{f}(x)_{RR}|f(x)=1)=\frac{(e^{\epsilon}+1)^{2}}{n(e^{\epsilon}-1)^{2}},$ (17)
which is always larger than the optimized $\widehat{MSE}$, i.e., $\frac{e^{\epsilon/2}}{n(e^{\epsilon}-1)^{2}}$. What is worse, even when $\epsilon$ is very large, it only converges to $\frac{1}{n}$, not zero. Therefore, before an analyst estimates the frequencies of a dataset, they need to determine whether to focus on (1) the overall accuracy (i.e., $l_{1}$ / $l_{2}$ losses) or (2) the worst-case accuracy of the individual estimations (i.e., $\widehat{MSE}$).
Since optimizing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses requires different $m$, we will hereinafter refer to the CMS optimized for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses as MSE-OCMS and $l$-OCMS, respectively. Additionally, the optimized CMS will be compared with the existing algorithms in Section 5.1.
@@ 3.5 Unbiased CMS with Imperfect Hashing
An optimized CMS is built upon an unbiased CMS, which requires its hashing to be pairwise independent. However, the construction of a $K$-wise-independent hashing family requires the range of the hashed values (i.e., $[m]$) to comprise a finite field [17]. A finite field must have $p^{l}$ elements [18], where $p$ is a prime number and $l\in\mathbb{N}^{+}$. For instance, if a CMS has $m=6$, there is no way to construct a pairwise independent hashing family for it. In practice, we can still construct an approximately pairwise-independent hashing family for any $m$. Consider a perfect pairwise independent hashing family with $p=2$ and $2^{l}\geq d$:
=$\mathcal{H}=\{h(x)|\forall a_{i}\in[2^{l}]:h(x)=a_{0}+a_{1}x\}$ (18)
where $a_{0}$, $a_{1}$, and $x$ are elements in a finite field $\mathbb{F}(2^{l})$, so the addition and multiplication in Eq. (18) are all field operations. The range of any $h(x)$ is also the field $\mathbb{F}(2^{l})$. Given two random variables $a_{0}$ and $a_{1}$, the size of $\mathcal{H}$ is $2^{2l}=\Theta(d^{2})$. Subsequently, we construct a new hashing family based on $\mathcal{H}$:
=$\mathcal{H}^{\prime}=\{h^{\prime}(x)|h(x)\bmod m\}.$
Since $\forall h\in\mathcal{H}$, $\forall y\in[m]$: $Pr(h(x)=y)=2^{-l}$, and assuming $2^{l}=qm+r$, where $q\in\mathbb{N}$ and $r\in\mathbb{N}$, we will have $r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q+1}{2^{l}}$ and the remaining $m-r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q}{2^{l}}$. Thus, we have:
@@
$Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=r(\frac{q+1}{2^{l}})^{2}+(m-r)(\frac{q}{2^{l}})^{2}\\ =\frac{(2q+1)r+mq^{2}}{(mq+r)^{2}},$ (19)
where $h^{\prime}\in\mathcal{H}^{\prime}$ represents uniform sampling.
Notice that $Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=\bar{c}(x_{1},x_{2})$, which is a constant and can be substituted into the $g(x)$ in Theorem 10, where $m^{\prime}=\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$. As a result, $g(x)$ is an unbiased estimator of $f(x)$. We can further prove that
@@
$Var(g(x|\mathcal{H}^{\prime}))<(1+\tau)Var(\hat{f}(x)),$
where $g(x|\mathcal{H}^{\prime})$ and $\hat{f}(x)$ represent the $g(x)$ with $\mathcal{H}^{\prime}$ and a standard estimator of $f(x)$ with a pairwise-independent hashing family.
Consider the optimized CMS for the worst-case MSE (Corollary 3.4), we can derive:
@@
When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon/2}$, we have
@@
Corollary 3.7.
When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon}$, we have

### v2

@@ 4. The frequency of any value $x\in[d]$ can be estimated as
Notably, $\hat{f}(x)$ may produce values outside the interval $[0,1]$ due to perturbations in the collected data. One straightforward solution is to truncate $\hat{f}(x)$ to $[0,1]$, i.e. $\min(1,\max(0,\hat{f}(x)))$, which also improves precision by eliminating deviations beyond the bounds. More advanced algorithms, such as projection onto a simplex, can also be used to enforce $\hat{f}(x)\in[0,1]$ and improve precision, as discussed in [6].
Regarding privacy guarantee, the only output that depends on the value of $X^{(i)}$ is $z^{(i)}$, which has been perturbed by an $\epsilon$-LDP mechanism. The postprocessing of the outputs from an LDP mechanism does not change its privacy guarantee [16], so we have proved:
@@ 3.1 Expectation and its Bias
$E[\hat{f}(x)]=\frac{m}{m-1}[f(x)\\ +\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1}$ (3)
@@
If and only if $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is satisfied for all $x\neq x^{\prime}$, then
@@
The proof of Theorem 3.2 and Corollary 3.1 is presented in Appendix D. It is notable that $\forall x\neq x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is equivalent to $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, where $h\in\mathcal{H}$ represents uniform sampling, which is also the definition of a pairwise-independent hashing family. Thus, Corollary 3.1 can be rephrased as:
@@ 3.2 Variance
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\sum_{x^{\prime}\in[d]}\biggl{[}\biggl{(}\frac{1}{k}\sum_{j\in[k]}Var\bigl{(}\\ R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\biggr{)}+\bar{c}(x^{\prime},x)-\bar{c}(x^{\prime},x)^{2}\biggr{]}f(x^{\prime}),$ (4)
@@
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\\ \biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq)+\\ \bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]},$ (5)
@@
The lemma, theorem, and corollary of this section are proven in Appendix E.
3.3 Accuracy and Precision Analysis
The mean squared error (MSE) measures the accuracy and precision of $\hat{f}(x)$, which is defined as
@@ 3.3 Accuracy and Precision Analysis
$MSE(g(x))=\frac{1}{n(1-\bar{c}(x,x^{\prime}))^{2}}\biggl{[}(1-f(x))\\ \biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq)\\ +\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (10)
@@
Besides MSE, one may also use the confidence interval (CI) of $\hat{f}(x)$ to measure precision. If $\hat{f}(x)$ is unbiased, its CI is the same as its concentration bound. Moreover, $\hat{f}(x)$ is equivalent to the sum of bounded random variables (i.e., $\sum\hat{y}$), so its bounds can be derived from the Chernoff bound. Below is the concentration bound for $\hat{f}(x)$ from the unbiased CMS using randomized response (referred to as CMS+RR):
@@
This theorem is proven in Appendix F. Notably, this theorem focuses solely on CMS+RR, as Section 3.4 will demonstrate that it outperforms CMS implementations with other known perturbations.
3.4 Optimizing CMS
This section explores the parameters of CMS to reduce the $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE defined in Section 2.2. It is notable that $l_{2}$ loss is connected to MSE as follows:
@@ 3.4 Optimizing CMS
We first consider randomized response (RR) as the LDP process, whose variance is as follows:
$Var(RR(v)[v^{\prime}])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}e^{\epsilon}(m-1)&\text{ if }v=v^{\prime}\\ e^{\epsilon}+m-2&\text{ if }v\neq v^{\prime}.\end{cases}$ (12)
Now, we optimize $m$ for CMS+RR, starting with the worst-case MSE. Here, we assume that an analyst has prior knowledge of the upper bound of $f(x)$ and derive:
@@
If $\forall x\in\mathbf{x}:f(x)\leq f^{*}$ is guaranteed to be correct by prior knowledge, where $\mathbf{x}$ are the values of interest, the worst-case MSE among $\mathbf{x}$ for CMS+RR is optimized as
@@
and the hash-function range $m$ is the closest integer to
@@
The theorem is proven in Appendix B. Notably, $f^{*}$ is prior knowledge of an analyst and could be larger than $\max_{x}f(x)$, which is unknown to the analyst. When $f^{*}\geq\frac{1}{2}$, it has the same worst-case MSE as $f^{*}=1$, which is guaranteed to be correct. On the other hand, if $f^{*}<\frac{1}{2}$ and the prior knowledge happens to be incorrect, i.e., $\max_{x}f(x)>f^{*}$, we demonstrate that
Theorem 3.7.
Given the optimized CMS+RR in Theorem 13, if $\delta=\max_{x}f(x)-f^{*}>0$ and $f^{*}<\frac{1}{2}$,
$\widehat{MSE}(\hat{f})=\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}}\\ \leq\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})}{n}.$ (14)
where $\widehat{MSE}(\hat{f})$ and $\widehat{MSE}(\hat{f}|f^{*})$ are the actual worst-case MSE and the worst-case MSE assuming $\max_{x}f(x)\leq f^{*}$, respectively.
The theorem is proven in Appendix B. Since $\widehat{MSE}(\hat{f}|f^{*})\geq\widehat{MSE}(\hat{f}|f^{*}=0)=\frac{4e^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$, as long as $\delta(1-2f^{*})=O(\frac{4e^{\epsilon}}{n(e^{\epsilon}-1)^{2}})$, we still have $\widehat{MSE}(\hat{f})=O(\widehat{MSE}(\hat{f}|f^{*}))$. Therefore, if the prior knowledge $f^{*}$ is guaranteed to be correct or its error is bounded by the above condition, considering the prior knowledge $f^{*}$ in the $\widehat{MSE}$ calculation is recommended. Section 6.2 provides an example of utilizing $f^{*}$. Otherwise, one can always assume $f^{*}=1$ for simplicity. Hereinafter, we assume no prior knowledge of $f(x)$ unless explicitly mentioned, i.e., $f^{*}=1$ by default.
Next, we optimize the $l_{1}$ and $l_{2}$ losses:
Theorem 3.8.
The $l_{2}$ loss and the upper bound of $l_{1}$ loss of CMS+RR are optimized as
$l_{2}^{*}(\hat{f})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}},$
and
respectively, where
$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)},$
and $m$ is the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$.
@@
When $d\gg e^{\epsilon}$, the optimized $l_{2}$ loss and the upper bound of the optimized $l_{1}$ loss of CMS+RR approach
$l_{2}^{*}(\hat{f})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$
@@
The theorem and corollary above are proven in Appendix B. One caveat is that the CMS+RR optimized for $l_{1}$ / $l_{2}$ losses is not optimized for worst-case MSE. Assuming $d$ is large enough, and $m=1+e^{\epsilon}$, we have
$\widehat{MSE}(\hat{f})=Var(\hat{f}(x)_{RR}|f(x)=1)=\frac{(e^{\epsilon}+1)^{2}}{n(e^{\epsilon}-1)^{2}},$
which is always larger than the optimized $\widehat{MSE}$ of CMS+RR, i.e., $\frac{e^{\epsilon/2}}{n(e^{\epsilon}-1)^{2}}$. What is worse, even when $\epsilon$ is very large, it only converges to $\frac{1}{n}$, not zero. Therefore, before an analyst estimates the frequencies of a dataset, they need to determine whether to focus on (1) the overall precision (i.e., $l_{1}$ / $l_{2}$ losses) or (2) the worst-case precision of the individual estimations (i.e., $\widehat{MSE}$).
Since optimizing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses requires different $m$, we will hereinafter refer to the CMS+RR optimized for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses as MSE-OCMS+RR and $l$-OCMS+RR, respectively. Additionally, OCMS+RR will be compared with the existing algorithms in Section 5.1.
Now, we evaluate the CMS using other existing LDP processes. A study [9] summarized the latest frequency estimation algorithms at that time. When considering different privacy regimes, they found that randomized response (RR), symmetric RAPPOR (RP), and asymmetric RAPPOR (a-RP) achieved state-of-the-art precision in some privacy regimes. Though more advanced algorithms exist, they are either as precise as these algorithms in their respective regimes, or they can be converted to CMS+RR (as shown in Section 4). Thus, in addition to RR, we will also evaluate CMS using RP and a-RP as the LDP process. Since the precision of these two mechanisms is independent of their input cardinality $m$, and $Var(\hat{f}(x))$ decreases with $m$ when $R$ is independent of $m$, the loss functions of CMS with RAPPOR mechanisms are optimized when $m\rightarrow\infty$. Appendix G derives their values as follows:
CMS+ RR $sRP$ $aRP$
$\widehat{MSE}(\hat{f}|$ $f^{*}<\frac{1}{2})$ $\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
$\widehat{MSE}(\hat{f}|$ $f^{*}\geq\frac{1}{2})$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
$l_{2}$ $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$ $\frac{de^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
CMS using RAPPOR mechanisms are less preferred for the following reasons: (1) their performance in reducing worst-case MSE or $l_{1}$ / $l_{2}$ losses is outperformed by CMS+RR in some conditions (highlighted in red in the table), and (2) they have very high communication costs (see Appendix G for details).
@@ 3.5 Unbiased CMS with Imperfect Hashing
OCMS+RR is built upon an unbiased CMS, which requires its hashing to be pairwise independent. However, the construction of a $K$-wise-independent hashing family requires the range of the hashed values (i.e., $[m]$) to comprise a finite field (Construction 3.32 of [17]). A finite field must have $p^{l}$ elements [18], where $p$ is a prime number and $l\in\mathbb{N}^{+}$. For instance, if a CMS has $m=6$, there is no way to construct a pairwise independent hashing family for it. In practice, we can still construct an approximately pairwise-independent hashing family for any $m$. Consider a perfect pairwise independent hashing family with $p^{l}>d$:
$\mathcal{H}_{pi}=\{h(x)=a_{0}+a_{1}x\mid a_{0},a_{1}\in[p^{l}]\}.$ (15)
where $a_{0}$, $a_{1}$, and $x$ are elements in a finite field $\mathbb{F}(p^{l})$, so the addition and multiplication in Eq. (15) are all field operations. The range of any $h(x)$ is also the field $\mathbb{F}(p^{l})$. One approach to construct $p^{l}$ is to find a prime number $p^{\prime}>d$ and set $l=1$, i.e., $p^{l}=p^{\prime}$. In this case, the field operations of $\mathbb{F}(p^{\prime})$ simply require every operation to modulo $p^{\prime}$ [19]. However, finding a prime number when $d$ is very large can be time-consuming. On the other hand, we can find any $l$ such that $2^{l}>d$, and set $p=2$, and the field operations of $\mathbb{F}(2^{l})$ are introduced in [19].
Subsequently, we construct a new hashing family based on $\mathcal{H}_{pi}$:
$\mathcal{H}_{api}=\{h^{\prime}(x)=h(x)\bmod m\mid h(x)\in\mathcal{H}_{pi}\}.$
where $api$ represents approximately pairwise independent. Notably, the finite field size $p^{l}$ should also be no less than $m$. Since $\forall h\in\mathcal{H}_{pi}$, $\forall y\in[m]$: $Pr(h(x)=y)=p^{-l}$, and assuming $p^{l}=qm+r$, where $q\in\mathbb{N}$ and $r\in\mathbb{N}$, we will have $r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q+1}{p^{l}}$ and the remaining $m-r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q}{p^{l}}$. Thus, we have:
@@
$Pr_{h^{\prime}\in\mathcal{H}_{api}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))\\ =r(\frac{q+1}{p^{l}})^{2}+(m-r)(\frac{q}{p^{l}})^{2}=\frac{(2q+1)r+mq^{2}}{(mq+r)^{2}},$ (16)
where $h^{\prime}\in\mathcal{H}_{api}$ represents uniform sampling.
Notice that $Pr_{h^{\prime}\in\mathcal{H}_{api}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=\bar{c}(x_{1},x_{2})$, which is a constant and can be substituted into the $g(x)$ in Theorem 10, where $m^{\prime}=\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$. As a result, $g(x)$ is an unbiased estimator of $f(x)$. We can further prove that
@@
$Var(g(x|\mathcal{H}_{api}))<(1+\tau)Var(\hat{f}(x)),$
where $g(x|\mathcal{H}_{api})$ and $\hat{f}(x)$ represent the $g(x)$ with $\mathcal{H}_{api}$ and a standard estimator of $f(x)$ with a pairwise-independent hashing family.
Applying the theorem to OCMS+RR, we can derive:
@@
When $2q+1>\sqrt{\frac{1}{\tau}}$, we have